# Contrasting exchange-field and spin-transfer torque driving mechanisms in all-electric electron spin resonance

### v1

@@ V Conclusions
In this work, we derived an intuitive physical picture of ESR in an STM, building on the solid mathematical aspects of non-equilibrium transport through a spin-polarized electrode under harmonic modulation of the bias. The key aspect of this model is the emergence of ESR driven by purely electric fields, which is experimentally realized in ESR-STM. We used a single-orbital Anderson impurity model with ionization and charging energies $\varepsilon,\varepsilon+U$, which is fairly general and can describe a wide variety of localized spin defects. Our model accounts for the driving amplitude in both the FLT and STT-driven regimes. However, it does not fully incorporate the complete transport picture, as cotunneling is not considered, leaving the door open for future extensions of this work.
@@
As we mentioned, the model has two distinctively different transport regimes. When the DC bias lies within the charging window $\varepsilon,\varepsilon+U$ the transport occurs in the Coulomb blockade (CB) regime. Here, the impurity is singly occupied and we found that ESR in this region is driven by an effective magnetic exchange field $B_{\mathrm{exch}}$ that arises as a consequence of the spin polarization of the electrode $P_{\alpha}$ i.e. the QI is driven by a FLT and we obtain results that aligns with conventional ESR following the Homodyne detection and presenting Rabi oscillations. Its component parallel to the spin quantization axis leads to a renormalization of the resonance frequency $f_{0}$ which is logarithmically dependent on the applied bias relative to the charging energy of the SAIM. This shift can be used to spectroscopically probe the impurity levels $\varepsilon,\varepsilon+U$ and the Rabi rate proportional to the exchange field. We further show that the quality factor, the product of Rabi rate and coherence time $\Omega_{\mathrm{FLT}}T_{2}$, has an optimum value close to the charging energy thresholds. In general, in the CB regime, the impurity spin is an ideal candidate for quantum-coherent control.
@@
When the bias crosses the charging thresholds, transport occurs outside the Coulomb blockade (CB) regime, allowing a full electron to hop in and out of the QI from the electrodes. Consequently, the current through the impurity increases, and $T_{2}$ decreases drastically. In this regime, an ESR signal can still be observed, with the driving mechanism now dominated by the spin polarized current, which torques the spin via the spin accumulation term and was referred as spin-transfer torque (STT). However, quantum-coherent control of the spin is not feasible due to rapid decoherence, as $\Omega_{\mathrm{STT}}T_{2}\sim 1/3$. Despite this, the QI can be strongly polarized along the spin’s quantization axis, provided it is not initialized perpendicular to the spin polarization. This offers an alternative method for initializing the spin in a preferred orientation, independent of temperature. Previously, STT was only studied using the semiclassical Landau–Lifshitz–Gilbert–Slonczewski equation that treats the magnetization as a classical vector. Thus, our model provides the first insight into quantum spin-transfer torque (qSTT).
@@
As a final remark, our model is able to reproduce a range of experimental observation in qualitative and quantitative good agreement, such as the homodyne detection and resonance shifts, which solidifies the validity of our driving mechanism approach and can serve as a guide for numerical optimization of on-surface electron spin systems for quantum simulations, quantum sensing and quantum information processing as well as a leap forward in the pursuit of a more complete theory able to encompass every transport regime i.e. both sequential and cotunneling regimes without discarding the exchange field from the sequential one.
@@ Acknowledgements.
We acknowledge fruitful discussions with A. Ardavan, A. J. Heinrich, C. Ast, C. P. Lutz, D.-J. Choi, D. Ruckert, D. Tomaszewski, F. Delgado, F. Donati, G. Burkard, G. Czap, H. T. Bui, J. Barnaś, J. König, J. C. Cuevas, L. Hoang-Anh, M. Kelai, M. Ternes, P. Kot, P. Busz, S.-H. Phark, S. Baumann, S. Loth, S. Stepanow, S. Kovarik and R. Schlitz during the process of writing this manuscript.
This work was supported by the Institute for Basic Science (IBS-R027-D1). Further financial support from projects RTI2018-097895-B-C44 and PID2021-127917NB-I00 funded by MCIN/AEI/10.13039/501100011033, QUAN-000021-01 funded by the Gipuzkoa Provincial Council and IT-1527-22 funded by the Basque Government is gratefully acknowledged. This work also received support from National Science Centre of Poland (Grant No. 2020/36/C/ST3/00539). Funded by the European Union. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union. Neither the European Union nor the granting authority can be held responsible for them.
@@ Appendix A Expectation value $\langle S_{i}\rangle$
A key consideration when computing the trace is ensuring that both operators are represented in the same basis. Otherwise, the expectation value would be incorrect. This is particularly relevant when considering an external magnetic field, as the quantization axis of the spin aligns with the magnetic field for an isotropic $g$-factor. Consequently, the density matrix and the Pauli matrices do not necessarily share the same basis, since the former is defined in the eigenstate basis, which is tilted by the magnetic field.
@@
where $l,j$ are singly occupied eigenstates of the spin-1/2, while $\sigma,\sigma^{\prime}$ correspond to the spin-up and spin-down states and $S_{i}$ are the Pauli matrices. We focus exclusively on the singly occupied states, as the empty and doubly occupied states provide zero spin to the simplest QI. For a general magnetic field, the singly eigenstates $l$ and $j$ have the following expression:
$|e\rangle$ $=$ $\left(\cos\frac{\theta}{2},e^{i\phi}\sin\frac{\theta}{2}\right),$ (32)
where $g$ and $e$ denote the ground and excite states of the QI respectively. Meanwhile, the standard up and down spin states basis set of the Pauli matrices are:
[…]
The density matrix is always expressed with the ground state as the first matrix element:
@@
For simplicity, we have been considering the case of $\phi=0$ for the derivation of Eqs. LABEL:spin_equations and throughout this work, arriving at the expressions for the spin of x and y used in the main text. From the spin components, we can obtain the population difference as well as the real and imaginary parts of the coherences, which are defined along the spin quantization axis:
@@ Appendix B Coefficients $\sum_{\sigma}\lambda_{lj,\sigma}\mu_{uv,\sigma}(1+2\sigma P_{\alpha}))$
These coefficients appear in the rates, and knowing how to evaluate them is essential for computing certain equations in this work. For the general eigenstates of a spin-1/2, Eqs. 32, the non-zero matrix elements between the ground/excited states and the transient ones are:
@@ Appendix C Details on current expression for a spin 1/2
$I_{L}(t)$ $=$ $\frac{2e}{\hbar}\left[\rho_{ee}(t)\mbox{Re}\left\{\Gamma_{e2,2e,L}^{+}(t)-\Gamma_{e0,0e,L}^{-}(t)\right\}\right.$ (34)
@@
$+$ $\sum_{j=e,g}\left(\rho_{0}(t)\mbox{Re}\left\{\Gamma_{0jj0,L}^{+}\right\}-\rho_{2}(t)\mbox{Re}\left\{\Gamma_{2jj2,L}^{-}\right\}\right)$
@@
where we have separated hole and electron contributions to the current and canceled out the ones that are zero due to being matrix elements $\lambda_{lj,\sigma}$ null. The major difference between the hole and electron rates comes from the first one containing $\varepsilon+U$, while the second one $\varepsilon$, but the spin dependence is similar and encoded in terms such as $\sum_{\sigma}\lambda_{lv\sigma}\lambda_{vj\sigma}^{*}(1+2\sigma P_{L})$, which we have already computed in the previous appendix. These expressions allow us to write:
$\Gamma_{ev,ve,L}^{\pm}(t)=\frac{1}{2}(1\mp P_{L}\cos\theta)\gamma_{L}(t){\cal I}_{L}^{\pm}(\Delta_{ve})$
@@
$\Gamma_{ev,vg,L}^{\pm}(t)=\mp\frac{1}{2}P_{L}\sin\theta\gamma_{L}(t){\cal I}_{L}^{\pm}(\Delta_{vg})$
$\sum_{j=e,g}\Gamma_{0j,j0,L}^{+}(t)=\gamma_{L}(t){\cal I}_{L}^{+}(\Delta_{j0})$
$\sum_{j=e,g}\Gamma_{2j,j2,L}^{-}(t)=\gamma_{L}(t){\cal I}_{L}^{-}(\Delta_{j2})$
where, on the first three equations, the ones where the first subscript in the rates start with a singly occupied state, the $+$ sign signifies the electron-like transport, implying that $v=0$, while the -- sign applies to the hole-like process, where $v=2$. Additionally, we have introduced the time dependent non-polarized coupling $\gamma_{L}(t)=\left(1+A_{L}\cos(\omega t)\right)^{2}\gamma_{L}^{0}$. Since the magnetic field is in the GHz regime, the integrals ${\cal I}_{L}^{\pm}(\Delta_{vj})$ will not depend on the Zeeman energy but only on $\varepsilon$, $U$, and $\mu_{\alpha}$ as we assumed in the main text, so we have:
@@
which together gives Eq. 12 when the population of the transient states, $\rho_{0}$ and $\rho_{2}$, are negligible and $\rho_{ee}(t)+\rho_{gg}(t)=\rho_{1}=1-\rho_{2}-\rho_{0}\approx 1$. We have employed Eq. 13. An analytical form of the current at low temperature is easy to obtain in terms of $\arctan$ functions.
Appendix D Insights on mapping the QME onto a spin problem
@@
Here we aim to clarify some of the steps done for mapping the spin 1/2 version of Eqs. 5 into a spin system obtaining Eqs. LABEL:spin_equations. This means that we will derive a set of equations that depends solely on the spin components and the transient populations of the QI. To begin with, let us write the full set of equations for the populations and coherences for the simplest QI:
$\hbar\dot{\rho}_{gg}$ $=$ $-2\sum_{v=0,2}\mbox{Re}(\Gamma_{gv,vg}(t))\rho_{gg}-\sum_{v=0,2}\Gamma_{ev,vg}(t)\rho_{ge}-\sum_{v=0,2}\Gamma_{ev,vg}^{*}(t)\rho_{eg}+2\sum_{v=0,2}\mbox{Re}(\Gamma_{vg,gv}(t))\rho_{v}$
@@
$\hbar\dot{\rho}_{ee}$ $=$ $-\sum_{v=0,2}\Gamma_{ev,vg}^{*}(t)\rho_{ge}-2\sum_{v=0,2}\mbox{Re}(\Gamma_{ev,ve}(t))\rho_{ee}-\sum_{v=0,2}\Gamma_{ev,vg}(t)\rho_{eg}+2\sum_{v=0,2}\mbox{Re}\Gamma_{ve,ev}(t)\rho_{v}$
$\hbar\dot{\rho}_{eg}$ $=$ $-\sum_{v=0,2}\Gamma_{ev,vg}^{*}(t)\rho_{gg}-\sum_{v=0,2}\Gamma_{ev,vg}(t)\rho_{ee}-\left[i\Delta_{ge}+\sum_{v=0,2}\left(\Gamma_{gv,vg}(t)+\Gamma_{ev,ve}^{*}(t)\right)\right]\rho_{eg}+2\sum_{v=0,2}\mbox{Re}(\Gamma_{vg,ev}(t))\rho_{v}$
$\hbar\dot{\rho}_{0}$ $=$ $2\mbox{Re}(\Gamma_{e0,0e}(t))\rho_{ee}+2\mbox{Re}(\Gamma_{g0,0g}(t))\rho_{gg}+2\mbox{Re}(\Gamma_{e0,0g}(t))(\rho_{ge}+\rho_{eg})-2\mbox{Re}(\Gamma_{0g,g0}(t)+\Gamma_{0e,e0}(t)))\rho_{0}$
$\hbar\dot{\rho}_{2}$ $=$ $2\mbox{Re}(\Gamma_{e2,2e}(t))\rho_{ee}+2\mbox{Re}(\Gamma_{g2,2g}(t))\rho_{gg}+2\mbox{Re}(\Gamma_{e2,2g}(t))(\rho_{ge}+\rho_{eg})-2\mbox{Re}(\Gamma_{2g,g2}(t)+\Gamma_{2e,e2}(t)))\rho_{2}.$
At this point, we need to make use of several relations for the rates that applies when both the magnetic field and the frequency of the time dependent field are small compared to the ionization and charge energies $\varepsilon$ and $\varepsilon+U$. Choosing $\pm$ depending on $v=0,2$ respectively we have that: $\mbox{Re}(\Gamma_{ev,ve})+\mbox{Re}(\Gamma_{gv,vg})=2\mbox{Re}(\Gamma_{ev,ve}^{P_{\alpha}=0}),$ $\mbox{Re}(\Gamma_{ev,ve})-\mbox{Re}(\Gamma_{gv,vg})=\pm 2\sum_{\alpha}P_{\alpha}\cos\theta\ \mbox{Re}(\Gamma_{gv,vg,\alpha}^{P_{\alpha}=0}),$ $\mbox{Im}(\Gamma_{ev,ve})+\mbox{Im}(\Gamma_{gv,vg})=2\mbox{Im}(\Gamma_{gv,vg}^{P_{\alpha}=0}),\mbox{Im}(\Gamma_{ev,ve})-\mbox{Im}(\Gamma_{gv,vg})=\pm 2\sum_{\alpha}P_{\alpha}\cos\theta\ \mbox{Im}(\Gamma_{ev,ve,\alpha}^{P_{\alpha}=0}),$ and $\Gamma_{ev,vg}=\Gamma_{gv,ve}$. In fact, in general, we can write: $\Gamma_{ev,ve}=\sum_{\alpha}\Gamma_{ev,ve,\alpha}^{P_{\alpha}=0}(1\pm P_{\alpha}\cos\theta)$ and $\Gamma_{gv,vg}=\sum_{\alpha}\Gamma_{gv,vg,\alpha}^{P_{\alpha}=0}(1\mp P_{\alpha}\cos\theta)$. With these relations and Eqs. 1314 and 15 we can write the spin system of equations:

### v2

@@ V Conclusions
In this work, we have derived an intuitive physical picture of ESR in an STM, building on the solid mathematical aspects of non-equilibrium transport through a spin-polarized electrode under harmonic modulation of the bias. Using a single-orbital Anderson impurity model (SAIM), which is fairly general and can describe a wide variety of localized spin defects, we capture the emergence of ESR driven purely by electric fields, which is experimentally realized in ESR-STM. Our model accounts for the driving amplitude in both the FLT-driven and the STT-driven regimes. However, it does not fully incorporate the complete transport picture, as cotunneling is not considered, leaving the door open for future extensions of this work.
@@
Two transport regimes emerge. Within the charging window ($\varepsilon,\varepsilon+U$), the transport occurs in the Coulomb blockade regime and the impurity is singly occupied, yielding ESR driven by the perpendicular component of the exchange field $B_{\mathrm{exch}}$ that arises as a consequence of the spin polarization of the electrode. This produces FLT-driven ESR, following results that align with conventional ESR, presenting homodyne detection and Rabi oscillations. On the other hand, the exchange field component parallel to the eigenbasis axis leads to a renormalization of the resonance frequency $f_{0}$, which is logarithmically dependent on the applied bias relative to the energy thresholds of the SAIM, $\varepsilon,\varepsilon+U$. This shift can be used to spectroscopically probe the energy thresholds and the Rabi rate. We further show that the quality factor $\Omega_{\mathrm{FLT}}T_{2}$ peaks near the charging energy thresholds. In general, in the Coulomb blockade regime, the impurity spin is an ideal candidate for quantum-coherent control.
@@
Outside the Coulomb blockade, transport allows electron hopping in and out of the quantum impurity from the electrodes. Consequently, the current through the impurity increases rapidly and $T_{2}$ decreases drastically. ESR is now STT-driven, dominated by the spin-polarized current, which torques the spin via the time-dependent spin accumulation, preventing coherent control ($\Omega_{\mathrm{STT}}T_{2}\sim 1/3$). Despite this, the quantum impurity can be strongly polarized along the spin’s eigenbasis axis, where the Hamiltonian is diagonal, provided it is not initialized perpendicular to the spin polarization. This offers an alternative method for initializing the spin in a preferred orientation, independently of temperature. Previously, STT was analyzed semiclassically using the Landau–Lifshitz–Gilbert–Slonczewski equation. Later, a quantum STT treatment explained the ESR signal in the cotunneling regime [31]. In contrast, our model captures quantum spin-transfer torque in sequential tunneling, including both coherent and incoherent ESR, highlighting the role of FLT-driven regime in spin control, a feature particularly relevant to ESR-STM.
@@
Finally, our model is able to reproduce a range of experimental observations in both qualitative and quantitative agreement, such as homodyne detection and resonance shifts. This solidifies the validity of our driving-mechanism approach and provides a guide for the numerical optimization of on-surface electron spin systems for quantum simulations, quantum sensing, and quantum information processing. Moreover, it constitutes a significant step toward a unified framework that treats both sequential and cotunneling transport regimes, while explicitly accounting for first-order coherent effects such as the exchange field or the Lamb shift.
@@ Acknowledgements.
We acknowledge fruitful discussions with A. Ardavan, A. J. Heinrich, C. Ast, C. P. Lutz, D.-J. Choi, D. Ruckert, D. Tomaszewski, F. Delgado, F. Donati, G. Burkard, G. Czap, H. T. Bui, J. Barnaś, J. König, J. C. Cuevas, L. Hoang-Anh, M. Kelai, M. Ternes, P. Gambardella, P. Kot, P. Busz, S.-H. Phark, S. Baumann, S. Loth, S. Stepanow, S. Kovarik, R. Broekhoven, and R. Schlitz during the process of writing this manuscript.
This work was supported by the Institute for Basic Science (IBS-R027-D1) and the German Research Foundation [Deutsche Forschungsgemeinschaft (DFG)] under Projects No. 450396347. Projects PID2021-127917NB-I00 funded by MCIN/AEI/10.13039/501100011033, QUAN-000021-01 funded by the Gipuzkoa Provincial Council, IT-1527-22 funded by the Basque Government, and ESiM project 101046364 supported by the European Union are gratefully acknowledged. Support from the National Science Centre of Poland (Grant No. 2020/36/C/ST3/00539) is also appreciated. Funded by the European Union. Views and opinions expressed are, however, those of the author(s) only and do not necessarily reflect those of the European Union. Neither the European Union nor the granting authority can be held responsible for them.
@@ Appendix A Expectation value $\langle s_{i}\rangle$
A key consideration when computing the trace is ensuring that both operators are represented in the same basis. Otherwise, the expectation value would be incorrect. This is particularly relevant when considering an external magnetic field, as the spin aligns with the magnetic field for an isotropic $g$-factor and the QI Hamiltonian will be diagonal in the axis parallel to $\mathbf{B}$. Consequently, the density matrix and the Pauli matrices do not necessarily share the same basis, since the former is defined in the eigenstate basis, which is tilted by the magnetic field.
@@
where $l,j$ are singly occupied eigenstates of the spin-1/2, while $\sigma,\sigma^{\prime}$ correspond to the spin-up and spin-down states and $s_{i}$ are the Pauli matrices. We focus exclusively on the singly occupied states, as the empty and doubly occupied states provide zero spin to spin-1/2 QI. For a general magnetic field, the singly ground and excited eigenstates in the up/down basis set have the following expression:
$|g\rangle=\left(\sin\frac{\theta}{2},-e^{i\phi}\cos\frac{\theta}{2}\right),\ |e\rangle=\left(\cos\frac{\theta}{2},e^{i\phi}\sin\frac{\theta}{2}\right),$ (33)
where $g$ and $e$ denote the ground and excite states of the QI respectively. Meanwhile, the standard up and down spin states of the Pauli matrices are: $|\mkern-5.0mu\downarrow\rangle=(0,1),\ |\mkern-5.0mu\uparrow\rangle=(1,0).$ The density matrix is always expressed with the ground state as the first matrix element:
@@
For simplicity, we have been considering the case of $\phi=0$ for the derivation of Eqs. (17a) and (17b) and throughout this work, arriving at the expressions for the spin of x and y used in the main text. From the spin components, we can obtain the population difference as well as the real and imaginary parts of the coherences, which are defined along the eigenbasis axis:
@@ Appendix B Coefficients $\sum_{\sigma}\lambda_{lj,\sigma}\mu_{uv,\sigma}(1+2\sigma P_{\alpha}))$
These coefficients appear in the rates, and knowing how to evaluate them is essential for computing certain equations in this work. For the general eigenstates of a spin-1/2, Eqs. (33),
$|e\rangle$ $=$ $\left(\cos\frac{\theta}{2},e^{i\phi}\sin\frac{\theta}{2}\right),$
the non-zero matrix elements between ground/excited states and the transient ones are:
@@ Appendix C Details on current expression for a spin 1/2
$I_{L}(t)$ $=$ $\frac{2e}{\hbar}\Bigg[\rho_{ee}(t)\mbox{Re}\left\{\Gamma_{e2,2e,L}^{+}(t)-\Gamma_{e0,0e,L}^{-}(t)\right\}$ (35)
@@
$+$ $\hskip-4.5pt\sum_{j=e,g}\left(\rho_{0}(t)\mbox{Re}\left\{\Gamma_{0jj0,L}^{+}\right\}-\rho_{2}(t)\mbox{Re}\left\{\Gamma_{2jj2,L}^{-}\right\}\right)\Bigg]$
@@
where we have separated hole and electron contributions to the current and canceled out the ones that are zero due to having matrix elements $\lambda_{lj,\sigma}$ null. The major difference between the hole and electron rates comes from the first one containing $\varepsilon+U$, while the second one $\varepsilon$, but the spin dependence is similar and encoded in terms such as $\sum_{\sigma}\lambda_{lv\sigma}\lambda_{vj\sigma}^{*}(1+2\sigma P_{L})$, which we have already computed in Appendix B. These expressions, taking into account that the magnetic field is in the GHz regime, so the integrals ${\cal I}_{L}^{\pm}(\Delta_{vj})$ will not depend on the Zeeman energy but only on $\varepsilon$, $U$, and $\mu_{\alpha}$, allow us to write:
$\Gamma_{ev,ve,L}^{\pm}(t)=\frac{1}{2}(1\mp P_{L}\cos\theta)\gamma_{L}(t){\cal I}_{L}^{\pm}(\Delta_{vg})$
@@
$\Gamma_{ev,vg,L}^{\pm}(t)=\mp\frac{1}{2}P_{L}\gamma_{L}(t){\cal I}_{L}^{\pm}(\Delta_{vg})\sin\theta$
$\sum_{j=e,g}\Gamma_{0j,j0,L}^{+}(t)=\gamma_{L}(t){\cal I}_{L}^{+}(\Delta_{g0})$
$\sum_{j=e,g}\Gamma_{2j,j2,L}^{-}(t)=\gamma_{L}(t){\cal I}_{L}^{-}(\Delta_{g2})$
where, on the first three equations, the ones where the first subscript in the rates start with a singly occupied state, the ++ sign signifies the electron-like transport, implying that $v=0$, while the - sign applies to the hole-like process, where $v=2$. Additionally, we introduce the time-dependent non-polarized coupling $\gamma_{L}(t)=\left[1+A_{L}\cos(\omega t)\right]^{2}\gamma_{L}^{0}$. Substituting these rates into Eq. (35) yields
@@
which together gives Eqs. (12) and (LABEL:homo_current_final), the latter by introducing Eq. (13) and when the population of the transient states, $\rho_{0}$ and $\rho_{2}$, are negligible and $\rho_{ee}(t)+\rho_{gg}(t)=\rho_{1}=1-\rho_{2}-\rho_{0}\approx 1$. For $E_{c}\gg|\varepsilon|,\varepsilon+U\gg\hbar/\tau_{c}$, ${\cal I}_{L}^{+}(\Delta_{e0})={\cal I}_{L}^{-}(\Delta_{e2})\rightarrow 0$ outside of the CB regime while $\rho_{0}=\rho_{2}=0$ inside of it, as Fig. 2c shows. This simplifies the equation of the current, obtaining the following analytical form at low temperatures:
$I_{L}(t)=\frac{e}{h}\gamma_{L}(t)$ (36)
$\times$ $\left[(\rho_{1}-2P_{L}\langle s_{z}\rangle)\left\{\frac{\pi}{2}+\arctan\left(\frac{eV_{\mathrm{DC}}-\varepsilon-U}{\hbar/\tau_{c}}\right)\right\}\right$.
- $\left.(\rho_{1}+2P_{L}\langle s_{z}\rangle)\left\{\frac{\pi}{2}-\arctan\left(\frac{eV_{\mathrm{DC}}-\varepsilon}{\hbar/\tau_{c}}\right)\right\}\right]$
Expression that gives zero if $\hbar/\tau_{c}=0$ and we are inside the CB regime, and
$I_{L}(t)=\pi e\ \mathrm{sign}(V_{\mathrm{DC}})\gamma_{L}(t)(\rho_{1}-2\ \mathrm{sign}(V_{\mathrm{DC}})P_{L}\langle s_{z}\rangle)/h,$
for DC bias values far above the energy thresholds, i.e. $|eV_{\mathrm{DC}}|\gg|\varepsilon|,\varepsilon+U$.
Appendix D Homodyne detection for a tilted magnetic-field plane
In this appendix we examine the dependence of the homodyne angle $\theta$ when the external magnetic field lies in a plane not containing the tip magnetization, as illustrated in Fig. 12. The plane is tilted by $\gamma=2$ degrees with respect to the spin-polarization direction. Increasing $\gamma$ gradually reduces the ESR signal until $\theta=90$ degrees, at which point homodyne detection vanishes and reemerges beyond this angle. This behavior arises because the ESR signal for in-plane orientations relative to the spin-polarization direction is STT-driven. Importantly, here $\theta$ denotes the angle between the tilted axis $z^{\prime}$ and the magnetic field, not between the tip magnetization (along $z$) and the field.
Figure 11: Schematic of the magnetic-field plane tilted by $\bm{\gamma}=2$ degrees (green), as shown in the zoom, formed between the tip’s magnetization direction, along $z$, and the new $z^{\prime}$ axis (in red). The magnetic field lies on the magenta curve with $\mathbf{B}=0.6$ T.
Figure 12: Homodyne detection for the tilted magnetic-field plane of Fig. 12. (a) $\theta$ dependence of the figure of merit $\Omega_{\mathrm{FLT}}T_{2}$ (black, left) and the total ESR intensity $\Delta I_{p}$ (blue, right). The overall behavior mirrors that of Fig. 10b and Fig. 9, but with a finite signal now observed at $\theta=0$ degrees and $180$ degrees, consistent with experiments. Parameters are identical to those of the main text at positive DC bias. (b) ESR intensity map, Eq. (32), versus homodyne angle and frequency, showing the ESR maximum near $\theta=10$ degrees, the flipping of the symmetric signal (peak-to-dip transition), and the evolution of the asymmetric component.
Figure 12a (blue) shows $\Delta I_{p}=\sqrt{\Delta I_{s}^{2}+\Delta I_{a}^{2}}$, where the small tilt of the magnetic field yields finite ESR signals at $\theta=0$ degrees and $180$ degrees, absent in the main-text results. This occurs because the magnetization direction of the tip can no longer be perfectly parallel to the magnetic field, a more realistic configuration in closer agreement with experiments [42]. For further discussion see Ref. [22], Fig. S15, noting that in experiments the tip is typically more in-plane, shifting the ESR maximum closer to $\theta=90$ degrees. For comparison, we also show $\Omega_{\mathrm{FLT}}T_{2}$ in Fig. 12a (in black), illustrating that its maximum does not coincide with the largest ESR signal.
Overall, the agreement between our model and experiment is remarkably good, again highlighting the importance of properly including the exchange field in homodyne detection. Without it, the ESR signal would resemble an STT-driven detection mechanism with a maximum at $\theta=90$ degrees, where the torque is strongest, which contradicts the experimental evidence, an issue we have already identified in other works [31, 36].
Figure 12b further presents a colormap of the raw ESR signal, Eq. (32), versus frequency and $\theta$. The plot reproduces the homodyne detection and reveals a sign reversal of the symmetric signal as $\theta$ approaches $90$ degrees from below, consistent with Figs. 5c and 6c. This flipping marks the crossover from FLT- to STT-driven detection, since the FLT contribution vanishes at $\theta=90$ degrees under homodyne detection, as discussed in the main text. The asymmetric component also evolves with angle, increasing from the out-of-plane configuration, peaking near $\theta=42$ degrees, and vanishing again for in-plane configurations, where the signal is STT-mediated. The exchange-field-induced frequency shift following Eq. (LABEL:energy_dress) is also visible in Fig. 12b.
Appendix E Insights on mapping the QME onto a spin problem
@@
Here we clarify the steps for mapping the spin-1/2 version of Eqs. (5) into a spin system, leading to Eqs. (17a) and (17b). We derive equations depending only on the spin components and transient QI populations. First, we write the full set for the populations and coherences for the simplest quantum impurity (spin 1/2):
$\hbar\dot{\rho}_{gg}$ $=$ $-2\sum_{v=0,2}\mbox{Re}(\Gamma_{gv,vg}(t))\rho_{gg}-\sum_{v=0,2}\Gamma_{ev,vg}(t)\rho_{ge}$
- $\sum_{v=0,2}\Gamma_{ev,vg}^{*}(t)\rho_{eg}+2\sum_{v=0,2}\mbox{Re}(\Gamma_{vg,gv}(t))\rho_{v}$
$\hbar\dot{\rho}_{ee}$ $=$ $-2\sum_{v=0,2}\mbox{Re}(\Gamma_{ev,ve}(t))\rho_{ee}-\sum_{v=0,2}\Gamma_{ev,vg}^{*}(t)\rho_{ge}$
- $\sum_{v=0,2}\Gamma_{ev,vg}(t)\rho_{eg}+2\sum_{v=0,2}\mbox{Re}\Gamma_{ve,ev}(t)\rho_{v}$
$\hbar\dot{\rho}_{0}$ $=$ $2\mbox{Re}(\Gamma_{e0,0e}(t))\rho_{ee}+2\mbox{Re}(\Gamma_{g0,0g}(t))\rho_{gg}$
$+$ $2\mbox{Re}(\Gamma_{e0,0g}(t))(\rho_{ge}+\rho_{eg})$
- $2\mbox{Re}(\Gamma_{0g,g0}(t)+\Gamma_{0e,e0}(t))\rho_{0}$
$\hbar\dot{\rho}_{2}$ $=$ $2\mbox{Re}(\Gamma_{e2,2e}(t))\rho_{ee}+2\mbox{Re}(\Gamma_{g2,2g}(t))\rho_{gg}$
$+$ $2\mbox{Re}(\Gamma_{e2,2g}(t))(\rho_{ge}+\rho_{eg})$
- $2\mbox{Re}(\Gamma_{2g,g2}(t)+\Gamma_{2e,e2}(t))\rho_{2}$
@@
$\hbar\dot{\rho}_{eg}$ $=$ $-\sum_{v=0,2}\Gamma_{ev,vg}^{*}(t)\rho_{gg}-\sum_{v=0,2}\Gamma_{ev,vg}(t)\rho_{ee}-\left[i\Delta_{ge}+\sum_{v=0,2}\left(\Gamma_{gv,vg}(t)+\Gamma_{ev,ve}^{*}(t)\right)\right]\rho_{eg}+2\sum_{v=0,2}\mbox{Re}(\Gamma_{vg,ev}(t))\rho_{v}.$
We need to make use of several relations for the rates that apply when both the magnetic field and the frequency of the time-dependent field are small compared to the energy thresholds $\varepsilon$ and $\varepsilon+U$. Choosing $\pm$ depending on $v=0,2$ respectively, we have:
$\mathrm{Re}(\Gamma_{ev,ve})+\mathrm{Re}(\Gamma_{gv,vg})$ $=$ $2\,\mathrm{Re}(\Gamma_{ev,ve}^{P_{\alpha}=0}),\ \ \mathrm{Re}(\Gamma_{ev,ve})-\mathrm{Re}(\Gamma_{gv,vg})=\pm 2\sum_{\alpha}P_{\alpha}\,\mathrm{Re}(\Gamma_{gv,vg,\alpha}^{P_{\alpha}=0})\cos\theta,$
$\mathrm{Im}(\Gamma_{ev,ve})+\mathrm{Im}(\Gamma_{gv,vg})$ $=$ $2\,\mathrm{Im}(\Gamma_{gv,vg}^{P_{\alpha}=0}),\ \ \mathrm{Im}(\Gamma_{ev,ve})-\mathrm{Im}(\Gamma_{gv,vg})=\pm 2\sum_{\alpha}P_{\alpha}\,\mathrm{Im}(\Gamma_{ev,ve,\alpha}^{P_{\alpha}=0})\cos\theta.$
and $\Gamma_{ev,vg}=\Gamma_{gv,ve}$. In fact, we can write $\Gamma_{ev,ve}=\sum_{\alpha}\Gamma_{ev,ve,\alpha}^{P_{\alpha}=0}(1\pm P_{\alpha}\cos\theta),\ \Gamma_{gv,vg}=\sum_{\alpha}\Gamma_{gv,vg,\alpha}^{P_{\alpha}=0}(1\mp P_{\alpha}\cos\theta)$. With these relations and Eqs. (13), (14), and (15) we write the spin system of equations: